BIMA ARRAY DETECTIONS OF HCN IN COMETS LINEAR (C/2002 T7) AND NEAT (C/2001 Q4)

\underline{D. N. FRIEDEL}, L. E. SNYDER, L. W. LOONEY, H. R. DICKEL, University of Illinois, Dept. of Astronomy, 1002 W. Green St, Urbana, IL 61801; ANTHONY J. REMIJAN, NASA Goddard Space Flight Center, Space and Earth Data Computing Division, Code 930, Greenbelt, MD 20771; National Research Council Resident Research Associate; PATRICK PALMER, Department of Astronomy and Astrophysics, University of Chicago, Chicago, IL 60637; IMKE DE PATER, J. R. FORSTER, C. KRAYBILL, M. C. H. WRIGHT, Department of Astronomy, University of California, Berkeley, CA 94720; M. R. HOGERHEIJDE, Leiden Observatory, PO Box 9513, 2300 RA, Leiden, The Netherlands; M. F. A’HEARN, Department of Astronomy, University of Maryland, College Park MD 20742-2421; and GEOFFREY A. BLAKE, Division of Geological and Planetary Sciences; Division of Chemistry and Chemical Engineering California Institute of Technology 150-21, Pasadena, CA 91125.

We present interferometric detections of HCN in comets LINEAR (C/2002 T7) and NEAT (C/2001 Q4) with the Berkeley-Illinois-Maryland Association (BIMA) Array in D-configuration. We detected the HCN J = 1 − 0 emission line in both comets. With a 25.′′4 × 20.′′3 synthesized beam around Comet LINEAR, we found a total beam averaged HCN column density of < N_{T} >= 8.5 ± 2.8 × 10^{12} cm^{−2}, and a HCN production rate of Q(HCN)=8.4 ± 2.8 × 10^{26} s^{−1}, giving a production rate ratio of HCN relative to H_{2}O of ∼ 1.6 ± 0.5 × 10^{−3} and relative to CN of ∼ 1.1 ± 0.4. With a 21.′′3 × 17.′′5 synthesized beam around Comet NEAT, we found a total beam averaged HCN column density of < N_{T} >= 1.0±0.5×10^{12} cm^{−2}, and a HCN production rate of Q(HCN)=9.2 ± 5.1 × 10^{25} s^{−1} giving a production rate ratio of HCN relative to H_{2}O of ∼ 0.8 ± 0.4 × 10^{−3} and relative to CN of ∼ 0.3 ± 0.2. For both comets, the production rates relative to H_{2}O are similar to previous comet observations. For Comet LINEAR the production rate relative to CN is consistent with HCN being the primary parent species of CN, while for Comet NEAT it is too low for this to be the case.